OpenStudy (anonymous):

2.Write a word problem describing the time it takes to complete an activity individually and together. (make up another character) For example, if John takes 2 hours to mow his lawn and it takes his sister, Maria, 4 hours to mow the same lawn, how long would it take John and Maria to mow the lawn together? 3.Write a rational equation based upon the word problem you created. .... any ideas for a question?

OpenStudy (anonymous):

time to type a term paper?

hero (hero):

Either look online for a similar problem or just make one up yourself using the example already given. There are plenty of online versions of solved "work problems". YouTube is a pretty good resource for this.

OpenStudy (anonymous):

If you call the two different times for completing the task x and y a general equation for solving all times can be written like this:\[x*y/(x+y)\] For the lawn mowing problem this gives us \[2*4/(2+4)=8/6=1.333...=1 hour 20 minutes\]